# A Survey on Knowledge Organization Systems of Research Fields: Resources and Challenges

### v1

@@ Abstract
Knowledge Organization Systems (KOSs), such as term lists, thesauri, taxonomies, and ontologies, play a fundamental role in categorising, managing, and retrieving information. In the academic domain, KOSs are often adopted for representing research areas and their relationships, primarily aiming to classify research articles, academic courses, patents, books, scientific venues, domain experts, grants, software, experiment materials, and several other relevant products and agents. These structured representations of research areas, widely embraced by many academic fields, have proven effective in empowering AI-based systems to i) enhance retrievability of relevant documents, ii) enable advanced analytic solutions to quantify the impact of academic research, and iii) analyse and forecast research dynamics. This paper aims to present a comprehensive survey of the current KOS for academic disciplines. We analysed and compared 45 KOSs according to five main dimensions: scope, structure, curation, usage, and links to other KOSs. Our results reveal a very heterogeneous scenario in terms of scope, scale, quality, and usage, highlighting the need for more integrated solutions for representing research knowledge across academic fields. We conclude by discussing the main challenges and the most promising future directions.
@@ 1 Introduction
In the academic domain, KOSs are often adopted for representing research areas and their relationships (Sugimoto and Lariviére, 2018), with the primary aim of classifying research articles, books, courses, patents, grants, software, experiment materials, scientific venues, domain experts, organisations, and several other relevant items and agents. These structured representation of knowledge have been adopted by most academic fields and proved very effective in i) improving the retrievability of relevant documents  (Newman et al., 2010; Salatino et al., 2019), ii) enabling advanced analytic solutions to quantify the impact of research (Ding et al., 2014; Qiu et al., 2017; Sugimoto and Lariviére, 2018; Salatino et al., 2023), and iii) understanding and forecasting research dynamics (Scharnhorst et al., 2012; Qiu et al., 2017).
@@
In line with the Open Science principles, we release the table that describes all the identified KOSs according to the 15 features on Open Research Knowledge Graph [^1] Full table describing the analysed KOSs according to the 15 features - https://doi.org/10.48366/R732033, as well as the code [^2] The code for processing the analysed KOSs is available on GitHub - https://github.com/angelosalatino/kos-rf we developed for processing them.
@@ 2.1 Knowledge Organization Systems in Digital Libraries
The annotation of documents based on the concepts within KOSs can be either performed manually or automatically. Manual annotation tasks are undertaken by human experts, typically experienced curators or editors, who leverage their domain knowledge to critically assess document content and assign the most pertinent concepts. In contrast, automatic annotation employs a range of computational tools that often incorporate advanced artificial intelligence techniques. For instance, OpenAlex, which is a major bibliographic catalogue of scientific papers, employs a deep learning model that, based on research papers’ title, abstract, citations, and journal name, can define the appropriate topics drawn from the OpenAlex Topics vocabulary (OpenAlex, 2024).
@@
In conclusion, digital libraries often employ KOS to improve document organisation and provide a wide range of advanced features.
@@ 2.2 Knowledge Organization Systems in Information Science
In the context of analysing the scientific landscape, Angioni et al. (2022a) developed the AIDA Dashboard, a tool that facilitates the analysis of conferences and journals in Computer Science, providing valuable insights into main authors, organisations, and countries. The dashboard leverages the Computer Science Ontology (Salatino et al., 2018b) to provide a very granular representation of the venues’ research topics, as well as to rank venues within topics using various metrics. Further contributions include Reymond (2020), who examined patents in the Humanities using the UNESCO Thesaurus. Their findings provided useful insights on potential research questions, and unexplored research avenues. Furthermore, the work of Yang and Lee (2018) introduced a tool that assists users in analysing research trends by expanding initial queries using MeSH terms, facilitating a more comprehensive exploration of the research landscape.
@@
These analyses and approaches underscore the value and importance of KOSs in research, as they provide the essential structure and organisation needed for developing various downstream applications.
@@ 3.1 Concepts
KOSs in this space primarily fall into four categories: term lists, hierarchical taxonomies, thesauri, and ontologies. In the following, we describe each one of them.
@@
A hierarchical taxonomy (from here on just taxonomy) organises the classes in a hierarchical structure with parent-child relationships (Rasch, 1987). In practical terms, as shown in Fig. 1(a), a taxonomy is typically organised in a tree structure, with a root node at the top that unfolds in several sub-branches. An important characteristic is that all items classified according to a class can also be considered under all its super-classes. For instance, in Fig. 1(a), all “Bacteria” are also “Organisms”.
@@
A thesaurus can be seen as an extension of the taxonomy. Subjects are organised in a hierarchical structure and they can also be described according to additional properties, such as an description, related terms, and synonyms (ANSI/NISO Z39.19-2005, R2010).
Finally, an ontology, from the “Information Technology” perspective, is a formal and explicit description of knowledge within a domain, categorising things according to their essential or relevant qualities (Gruber, 1993). Ontologies conceptualise a subject area according to an abstract model of how domain experts rationalise knowledge in the domain. In practice, ontologies consist of a set of concepts, objects, and other entities and the relationships between them (Genesereth and Nilsson, 2012). Compared to thesauri, ontologies are more expressive since they can also express axioms and restrictions, so to provide local constraints on properties. As an example, Fig. 1(b) shows a portion of the MeSH ontology. In particular, the two mesh:broaderDescriptor relationships define the super-class of both “Alveolata” and “Amoebozoa” which is “Eukaryota”. Next, rdfs:label provides a human-readable version of a given class. Moreover, the mesh:previousIndexing points to the same element in a previous version of the KOS, and the mesh:annotation provides a human readable description of the class.
@@ 3.3 Methodology for the retrieval of relevant KOSs
Identifying all the relevant KOSs has proven to be a challenging task. A common approach, when pursuing a survey or a review of the literature, is to rely on a search engine, e.g., Scopus, Web of Science, Google Scholar, and construct a particular query that would return all research papers reporting the objects under review (Pranckutė, 2021). However, a good number of systems organising research areas are not well described or documented in research papers. Therefore, relying on this typical approach would have produced limited results. To this end, we designed and performed the following systematic strategy.
  1. Phase 1:
    We started by querying Google Scholar for potential candidates using the following query “(controlled vocabulary OR taxonomy OR thesaurus OR ontology OR subject headings OR subject classification) AND (research OR science)”.
    Due to Google Scholar’s limitations with logical operators, we executed the queries individually and then combined the resulting papers. In this phase, we gathered: i) WikiCSSH(Han et al., 2020), ii) Computer Science Ontology(Salatino et al., 2018b), iii) Medical Subject Headings(Darmoni et al., 2012), iv) Library of Congress Classification(Summers et al., 2008), v) Modern Science Ontology(Fathalla et al., 2023), vi) Science Metrix Classification(Archambault et al., 2011), and vii) OpenAIRE’s Field of Science Taxonomy(Kotitsas et al., 2023).
  2. Phase 2:
    We analysed the websites of academic publishers, pre-print archives, and academic search engines to identify the KOSs they use to organise their content. Specifically, we considered Scopus, Web of Science, Microsoft Academic Graph, Dimensions.ai, PubMed, the Springer Nature portal, ACM, IEEE, OpenAlex, and ArXiv. As a result, we identified: i) ACM Computing Classification Scheme, ANZSRC ii) Fields of Research, iii) ArXiv Subjects, Microsoft’s, iv) Fields of Study, v) Nature Subjects, vi) IEEE Thesaurus, vii) Web of Science Categories, viii) All Science Journal Classification Codesof Scopus, and ix) OpenAlex Topics.
  3. Phase 3:
    We adopted the Google search engine to identify additional KOSs, using the query “("controlled vocabulary" OR "term list" OR "taxonomy" OR "thesaurus" OR "ontology" OR "subject headings" OR "subject classification")” in combination with the 19 broad fields listed in IC 2. As a result, we identified: i) Art and Architecture Thesaurus, ii) Biomedical Ontologies from BioPortal, iii) ChemOnt, iv) GeoRef Thesaurus, v) KNOWMAK, vi) Journal of Economic Literature, vii) Mathematics Subject Classification, viii) U.S. Geological Survey Library Classification System, ix) PhilPapers Taxonomy, x) Physical Subject Headings, xi) PsycInfo and PsycTests Classification Systems, and xii) Subject Resource Application Ontology.
  4. Phase 4:
    We contacted researchers working in the field of “Digital Libraries” and asked them whether they could point us toward any additional effort. As a result, we gathered: i) Dewey Decimal Classificationand ii) Unified Medical Language System.
  5. Phase 5:
    We expanded the resulting KOSs by analysing their links on Wikipedia. Specifically, we relied on the “See also” section, which typically lists online databases and other related KOSs. As a result, we gathered: i) STW Thesaurus for Economics, ii) Physics and Astronomy Classification Scheme, iii) Open Biological and Biomedical Ontology, iv) National Library of Medicine classification, and v) Fields of research and development.
  6. Phase 6:
    We retrieved and considered all the KOSs that are explicitly linked into the already selected ones. Thanks to this process, we included: i) Agrovoc Thesaurus, ii) DFG Classification, iii) EDAM, iv) TheSoz, ANZSRC’s v) Socio-Economic Objective, ASRC’s vi) Research Fields, Courses and Disciplines, vii) EuroVoc, and viii) Unesco Thesaurus.
  7. Phase 7:
    Whenever we could not find at least one KOS within one of the 19 disciplines defined in IC2 (e.g., “History”), we reached out to professors in the missing disciplines, asking which KOS they employ, if any. Their response directed us toward multi-field KOS such as i) European Commission Taxonomyand ii) European Research Council Taxonomy.
@@ 3. Semi-automatic
Figure 2: The features used for the analysis of knowledge organization systems.
Scope. The scope is the set of research fields covered by a KOS. Some KOSs focus on one specific field, such as PhySH for “Physics” or Mathematics Subject Classification for “Mathematics”. These may also include elements from other complementary fields. This is the case of PhilPapers Taxonomy that mainly focuses on “Philosophy”, but also describes some concepts from other fields, including “Mathematics”. Another category of KOSs covers multiple fields by design, typically because they aim to offer good coverage of the full set of scientific or academic disciplines. A good example is the Unesco Thesaurus, a well-known knowledge organization system which aims to cover all academic disciplines and is adopted by several libraries.
@@
We then considered the number of concepts and the maximum depth of the hierarchical tree (which may be one in the case of term lists). The depth was computed as the number of levels from the most generic concept (root) to the most specific concepts (leaves). Depth can be generally used as an indicator of specificity, since deep taxonomies tend to include a granular representation of very detailed (narrow) topics and may allow for more fine-grained content organisation.
@@
We used Python scripts for automatically extracting the structural features from the KOSs published in standard formats (e.g., RDF). The scripts are publicly available as Jupyter notebooks on a GitHub repository [^10] Jupyter notebooks for analysing KOSs - https://github.com/angelosalatino/kos-rf. For the knowledge organization systems that are available only as HTML or PDF, we manually extracted these features.
Curation. The curatorial aspect takes into consideration many features regarding the creation, update, and publication of the KOS. First, we considered how the KOS was generated according to four categories: manual, automatic, semi-automatic, and community effort. The first and more common category includes all the KOSs that were crafted manually by domain experts, such as the ACM Computing Classification Scheme. The automatic category covers the KOSs that were generated automatically using NLP or Machine Learning approaches, such as the Computer Science Ontology. The semi-automatic category covers KOSs generated semi-automatically. For instance, the OpenAlex Topics uses a manual approach for the first three levels and an automatic one for the last level. The community effort category includes the KOSs that were crowdsourced with the support of the community, such as Physical Subject Headings.
We also considered the format used for publishing the KOSs. A good number of KOSs are published according to W3C standard formats such as Ontology Web Language (OWL), Terse RDF Triple Language (TTL), N-Triples (NT) for expressing data in the Resource Description Framework (RDF) [^11] Resource Description Framework (RDF) — https://www.w3.org/RDF data model, as well as JSON, and CSV. An alternative is the OBO format [^12] OBO Format — http://purl.obolibrary.org/obo/oboformat/spec.html, which is still a working draft standard for representing ontologies, based on the principles of OWL and predominantly used in the field of “Biology”. Other KOSs are instead published in semi-structured formats, for instance, as a list of terms in HTML pages or PDF files. These solutions are typically not machine-readable and offer very limited support to automatic classifiers.
@@
External Links. Links to other KOSs allow the integration of different knowledge bases, potentially offering a more comprehensive representation of scientific disciplines. Therefore, we documented whether a KOS includes links to other KOSs, including those referencing general knowledge graphs, such as DBpedia [^14] DBpedia — https://wiki.dbpedia.org and Wikidata [^15] Wikidata — https://www.wikidata.org. For example, the STW Thesaurus for Economics is mapped to Wikidata, DBpedia, the JEL classification, and others. We will discuss the crucial implications of this mechanism in Sec. 5.
Usage. The final aspect that we considered regards the presence of collections of documents or other artefacts that are annotated or organised according to the KOS. Their existence attests to the fact that a portion of the community actively uses KOS. Furthermore, annotated datasets can also be used for training machine learning classifiers able to categorise new documents according to the original KOS (Kandimalla et al., 2021; Salatino et al., 2021). For instance, the Fields of Research is currently adopted by institutional repositories in both Australia and New Zealand as well as Dimensions.ai [^16] Dimension.ai — https://app.dimensions.ai/discover/publication to index their publications metadata. Another example is Physical Subject Headings employed to classify the articles of the Physical Review Journal [^17] Physical Review Journal — https://journals.aps.org.
@@ 4 Analysis of Knowledge Organization Systems
This section analyses the KOSs of research fields utilising the 15 features reported in Fig. 2. We identified 45 KOSs: 23 specialising in single disciplines and 22 covering multiple fields. Table 1 summarises these KOSs according to some key characteristics, i.e., primary discipline, number of concepts, hierarchical depth, type of system, kind of hierarchy, and whether it contains related terms. An extended version of this table with all the analysed features is available at https://doi.org/10.48366/R732033.
The following subsections discuss each of the five aspects outlined in Section 3.4, highlighting trends and providing specific examples.
@@ 4.1 Scope
Out of the 45 identified KOS, 22 cover multiple fields, while 23 focus on a single field of science. However, only 12 of the 19 fields under analysis are addressed by at least one single-field KOS, while the remaining ones rely exclusively on broader multi-field solutions.
Notably, 6 fields out of the 19 under analysis (see IC 2, Sec. 3.2) are covered by more than one KOS. In “Medicine”, we identified four single-field KOSs: the Medical Subject Headings, the Unified Medical Language System, the National Library of Medicine classification, and the Biomedical Ontologies from BioPortal. Although the last expands toward “Biomedicine” and hence more into the biological basis of health and disease.“Computer Science” can rely on three KOSs: the Computer Science Ontology, the ACM Computing Classification Scheme, and Computer Science Subject Headings from Wikipedia. In the field of “Biology”, we identified two KOSs: the Open Biological and Biomedical Ontology and EDAM. However, EDAM focuses more on “Bioinformatics”, describing the applied side of “Biology” with “Computer Science” tools. In the field of “Geology”, we found the GeoRef Thesaurus and the U.S. Geological Survey Library Classification System. “Economics” is also covered by two KOSs: the Journal of Economic Literature Classification System and the STW Thesaurus for Economics. “Physics” can also rely on the Physics and Astronomy Classification Scheme and Physical Subject Headings.
Most other fields rely on just one KOS. This is the case for “Agriculture” (Agrovoc Thesaurus), “Art and Architecture” (Art and Architecture Thesaurus), “Chemistry” (ChemOnt), “Engineering” (IEEE Thesaurus), “Mathematics” (Mathematics Subject Classification), “Philosophy” (PhilPapers Taxonomy), and “Psychology” (PsycInfo and PsycTests Classification Systems).
Four KOSs were not directly assigned to one of the standard 19 broad fields, as they focus on more specialised areas. Specifically, we associated Agrovoc to “Agriculture”, which is a sub-area of the “Environmental Science”; Biomedical Ontologies from BioPortal to “Biomedicine”, which covers both “Biology” and “Medicine”; EDAM to “Bioinformatics”, which includes both “Biology” and “Computer Science”; and TheSoz to “Social Science”, which is typically considered a super-area of “Economics”, “Sociology”, “Psychology”, and “Political Science”.
We were unable to identify a single-field KOS for the following seven fields: “History”, “Political Science”, “Environmental Science”, “Material Science”, “Geography”, “Sociology”, and “Business”. We reached out to a number of professors and domain experts in such fields, who confirmed this finding and mentioned that they usually rely on generic multi-field KOSs such as the Dewey Decimal Classification and the Library of Congress Classification. To the best of their knowledge, KOSs for such fields are yet to be developed.

### v2

@@ Abstract
Knowledge Organization Systems (KOSs), such as term lists, thesauri, taxonomies, and ontologies, play a fundamental role in categorising, managing, and retrieving information. In the academic domain, KOSs are often adopted for representing research areas and their relationships, primarily aiming to classify research articles, academic courses, patents, books, scientific venues, domain experts, grants, software, experiment materials, and several other relevant products and agents. These structured representations of research areas, widely embraced by many academic fields, have proven effective in empowering AI-based systems to i) enhance the retrievability of relevant documents, ii) enable advanced analytic solutions to quantify the impact of academic research, and iii) analyse and forecast research dynamics. This paper aims to present a comprehensive survey of the current KOS for academic disciplines. We analysed and compared 45 KOSs according to five main dimensions: scope, structure, curation, usage, and links to other KOSs. Our results reveal a very heterogeneous scenario in terms of scope, scale, quality, and usage, highlighting the need for more integrated solutions for representing research knowledge across academic fields. We conclude by discussing the main challenges and the most promising future directions.
@@ 1 Introduction
In the academic domain, KOSs are often adopted for representing research areas and their relationships (Sugimoto and Larivière, 2018), with the primary aim of classifying research articles, books, courses, patents, grants, software, experiment materials, scientific venues, domain experts, organisations, and several other relevant items and agents. These structured representation of knowledge have been adopted by most academic fields and proved very effective in i) improving the retrievability of relevant documents  (Newman et al., 2010; Salatino et al., 2019), ii) enabling advanced analytic solutions to quantify the impact of research (Ding et al., 2014; Qiu et al., 2017; Sugimoto and Larivière, 2018; Salatino et al., 2023), and iii) understanding and forecasting research dynamics (Scharnhorst et al., 2012; Qiu et al., 2017).
@@
In adherence to Open Science principles, we release the table that describes all the identified KOSs according to the 15 features on Open Research Knowledge Graph [^1] Full table describing the analysed KOSs according to the 15 features — https://doi.org/10.48366/R732033, as well as the code [^2] The code for processing the analysed KOSs is available on GitHub — https://github.com/angelosalatino/kos-rf we developed for processing them.
@@ 2.1 Knowledge Organization Systems in Digital Libraries
The annotation of documents based on the concepts within KOSs can be either performed manually or automatically. Manual annotation tasks are undertaken by human experts, typically experienced curators or editors, who leverage their domain knowledge to critically assess document content and assign the most pertinent concepts. In contrast, automatic annotation employs a range of computational tools that often incorporate advanced artificial intelligence techniques (Salatino et al., 2019). For instance, OpenAlex, which is a major bibliographic catalogue of scientific papers, employs a deep learning model that, based on research papers’ title, abstract, citations, and journal name, can define the appropriate topics drawn from the OpenAlex Topics vocabulary (OpenAlex, 2024).
@@
In conclusion, digital libraries often employ KOSs to improve document organisation and provide a wide range of advanced features.
@@ 2.2 Knowledge Organization Systems in Information Science
In the context of analysing the scientific landscape, Angioni et al. (2022a) developed the AIDA Dashboard, a tool that facilitates the analysis of conferences and journals in Computer Science, providing valuable insights into main authors, organisations, and countries. The dashboard leverages the Computer Science Ontology (Salatino et al., 2018b) to provide a very granular representation of the venues’ research topics, as well as to rank venues within topics using various metrics. Further contributions include Reymond (2020), who examined patents in the Humanities using the UNESCO Thesaurus. Their findings provided useful insights into potential research questions and unexplored research avenues. Furthermore, the work of Yang and Lee (2018) introduced a tool that assists users in analysing research trends by expanding initial queries using MeSH terms, facilitating a more comprehensive exploration of the research landscape.
@@
These analyses and approaches underscore the value and importance of KOSs in research, as they provide the essential structure and organisation needed for developing various downstream applications (Salatino et al., 2020a).
@@ 3.1 Concepts
KOSs in this space primarily fall into four possible types of system: term lists, hierarchical taxonomies, thesauri, and ontologies. In the following, we describe each one of them.
@@
A hierarchical taxonomy (from here on just taxonomy) organises the classes in a hierarchical structure with parent-child relationships (Rasch, 1987). In practical terms, as shown in Figure 1(a), a taxonomy is typically organised in a tree structure, with a root node at the top that unfolds in several sub-branches. An important characteristic is that all items classified according to a class can also be considered under all its super-classes. For instance, in Figure 1(a), all “Bacteria” are also “Organisms”.
@@
A thesaurus can be seen as an extension of the taxonomy. Subjects are organised in a hierarchical structure and they can also be described according to additional properties, such as a description, related terms, and synonyms (ANSI/NISO Z39.19-2005, R2010).
Finally, an ontology, from the “Information Technology” perspective, is a formal and explicit description of knowledge within a domain, categorising things according to their essential or relevant qualities (Gruber, 1993). Ontologies conceptualise a subject area according to an abstract model of how domain experts rationalise knowledge in the domain. In practice, ontologies consist of a set of concepts, objects, and other entities and the relationships between them (Genesereth and Nilsson, 2012). Compared to thesauri, ontologies are more expressive since they can also express axioms and restrictions, so to provide local constraints on properties. As an example, Figure 1(b) shows a portion of the MeSH ontology. In particular, the two mesh:broaderDescriptor relationships define the super-class of both “Alveolata” and “Amoebozoa” which is “Eukaryota”. Next, rdfs:label provides a human-readable version of a given class. Moreover, the mesh:previousIndexing points to the same element in a previous version of the KOS, and the mesh:annotation provides a human-readable description of the class.
@@ 3. Semi-automatic
Figure 2: The aspects and features used for the analysis of Knowledge Organization Systems.
Scope. The scope is the set of research fields covered by a KOS. Some KOSs focus on one specific field, such as PhySH for “Physics” or Mathematics Subject Classification for “Mathematics”. These may also include elements from other complementary fields. This is the case of PhilPapers Taxonomy that mainly focuses on “Philosophy”, but also describes some concepts from other fields, including “Mathematics”. Another category of KOSs covers multiple fields by design, typically because they aim to offer good coverage of an extended set of scientific or academic disciplines. A good example is the UNESCO Thesaurus, a well-known KOS which aims to cover all academic disciplines and is adopted by several libraries.
@@
We then considered the number of concepts and the maximum depth of the hierarchical tree (which is one in the case of term lists). The depth was computed as the number of levels from the most generic concept (root) to the most specific concepts (leaves). Depth can be generally used as an indicator of specificity, since deep taxonomies tend to include a granular representation of very detailed (narrow) topics and may allow for more fine-grained content organisation.
@@
We used Python scripts for automatically extracting the structural features from the KOSs published in standard formats (e.g., RDF). The scripts are publicly available as Jupyter notebooks on a GitHub repository [^10] Jupyter notebooks for analysing KOSs — https://github.com/angelosalatino/kos-rf. For the knowledge organization systems that are available only as HTML or PDF, we manually extracted these features.
Curation. The curatorial aspect takes into consideration many features regarding the creation, update, and publication of the KOS. First, we considered how the KOS was generated according to four categories: manual, automatic, semi-automatic, and community effort. The first and more common category includes all the KOSs that were crafted manually by domain experts, such as the ACM Computing Classification Scheme. The automatic category covers the KOSs that were generated automatically using NLP or Machine Learning approaches, such as the Computer Science Ontology. The semi-automatic category covers KOSs generated by blending domain experts and automatic tools. For instance, the OpenAlex Topics uses a manual approach for the first three levels and an automatic one for the last level. The community effort category includes the KOSs that were crowdsourced with the support of the community, such as Physical Subject Headings.
We also considered the format used for publishing the KOSs. A good number of KOSs are published according to W3C standard formats such as Ontology Web Language (OWL), Terse RDF Triple Language (TTL), N-Triples (NT) for expressing data in the Resource Description Framework (RDF) [^11] Resource Description Framework (RDF) — https://www.w3.org/RDF data model, as well as JSON, and CSV. An alternative is the OBO format [^12] OBO Format — http://purl.obolibrary.org/obo/oboformat/spec.html, which is still a working draft standard for representing ontologies, based on the principles of OWL and predominantly used in the field of “Biology”. Other KOSs are instead published in semi-structured formats, for instance, as a list of terms in HTML pages or PDF files. These solutions are typically not machine-interpretable and offer very limited support to automatic classifiers.
@@
External Links. Links to other KOSs allow the integration of different knowledge bases, potentially offering a more comprehensive representation of scientific disciplines. Therefore, we documented whether a KOS includes links to other KOSs, including those referencing general knowledge graphs, such as DBpedia [^14] DBpedia — https://wiki.dbpedia.org and Wikidata [^15] Wikidata — https://www.wikidata.org. For example, the STW Thesaurus for Economics is mapped to Wikidata, DBpedia, and others. We discuss the crucial implications of this mechanism in Section 5.
Usage. The final aspect that we considered regards the presence of collections of documents or other artefacts that are annotated or organised according to the KOS. Their existence attests to the fact that a portion of the community actively uses that KOS. Furthermore, annotated datasets can also be used for training machine learning classifiers able to categorise new documents according to the original KOS (Kandimalla et al., 2021; Salatino et al., 2021). For instance, the Fields of Research is currently adopted by institutional repositories in both Australia and New Zealand as well as Dimensions.ai [^16] Dimension.ai — https://app.dimensions.ai/discover/publication to index their publications metadata. Another example is Physical Subject Headings employed to classify the articles of the Physical Review Journal [^17] Physical Review Journal — https://journals.aps.org.
@@ 4 Analysis of Knowledge Organization Systems
This section analyses the KOSs of research fields utilising the 15 features reported in Figure 2. We identified 45 KOSs: 23 specialising in single disciplines and 22 covering multiple fields. Table 1 summarises these KOSs according to some key characteristics, i.e., primary discipline, number of concepts, hierarchical depth, type of system, kind of hierarchy, and whether it contains related terms. An extended version of this table with all the analysed features is available at https://doi.org/10.48366/R732033.